OpenStudy (anonymous):

A particle moves at a velocity given by v(t)=6sin(t/2). If the particle is at x=1 when t=0, then its position when t=pi is x=?

OpenStudy (anonymous):

do you know what the position function would be?

OpenStudy (anonymous):

I think it would be the antiderivative?

OpenStudy (isaiah.feynman):

Nope, it will be the integral of the velocity function. @Euler271 thanks for mentioning position function I forgot it at first and was pondering where does x come from?

OpenStudy (anonymous):

The integral of the velocity function?

OpenStudy (anonymous):

\[\int\limits_{a}^{b}v(t)dt\]

OpenStudy (anonymous):

Something like that?

OpenStudy (anonymous):

integral = anti-derivative that is correct

OpenStudy (anonymous):

tell me the indefinite integral of v(t) [no boundaries]

OpenStudy (anonymous):

\[\int\limits_{?}^{?}6\sin(\frac{ t }{ 2 })\]

OpenStudy (anonymous):

not quite. let u = t/2 du = dt/2 you get \[\int\limits_{}^{}\sin \left(\frac{ t }{ 2 }\right)= -12\cos \left(\frac{ t }{ 2 }\right) + C\] the + C needs to be added when you take the indefinite integral (meaning no boundaries)

OpenStudy (anonymous):

i meant 6sin(t/2)dt

OpenStudy (anonymous):

Oh. I wasn't quite sure what you meant. However, I did get that equation, too.

OpenStudy (anonymous):

to find what C is for this particular scenario, you need to plug in the given value x(t) = -12cos(t/2) + C

OpenStudy (anonymous):

when t = 0, x(t) = 1 solve for C

OpenStudy (anonymous):

sorry for long time to reply

OpenStudy (isaiah.feynman):

So in essence its just 1=-12+C

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That's what I got.

OpenStudy (anonymous):

so now x(t) = -12cos(t/2) + 13 you can now solve for x(t) for any given t. where x(t) is the position x at time t

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

glad to help :)